Gary runs 1 km/h faster than Margie.
They run in opposite directions for 2 hours.
They are then 30 km apart.
How many km/h does Margie run?

Answer by fractalier(2101)

Remember D = RT, distance = rate*time.
Let Margie's rate be r.
Then Gary's rate is r + 1.
The sum of their distances is 30 km.
So, since they ran for 2 hours, t = 2, and we have
2r + 2(r + 1) = 30
2r + 2r + 2 = 30
4r + 2 = 30
4r = 28
r = 7, Margie's rate.